Calculate the ratio of the critical angle of diamond to the critical angle of crown glass if both are surrounded by a vacuum.


1
Expert's answer
2021-11-14T17:14:12-0500

The critical angle for a material is


"\theta=\sin^{-1}\frac 1n."

Therefore, for the diamond and crown glass, we have


"\frac{\theta_d}{\theta_g}=\frac{\arcsin(1/n_d)}{\arcsin(1/n_g)}=\frac{\arcsin(1/2.417)}{\arcsin(1/1.517)}=0.59."